# Deriving a GENERIC system from a Hamiltonian system

### v2

@@ 2.4 Invariant measure on the product space ${\mathbf{Z}}{\times}{\mathbf{X}}$
$\exp\left(-\beta{\mathcal{H}}_{\mathrm{total}}(z,\eta)\right)=\exp\left(-\beta\big{(}{\mathcal{H}}_{\mathcal{A}}(z)+\frac{1}{2}\|{\mathsf{C}}z\|_{\mathbf{H}}^{2}\big{)}\right)\exp\left(-\frac{\beta}{2}\|\eta{+}{\mathsf{C}}z\|_{\mathbf{H}}^{2}\right).$
@@
By introducing the new variables $(z,\zeta)=\widetilde{M}_{t}(z,\eta)=\big{(}z,\widehat{S}_{-t}(\eta{+}{\mathsf{C}}z)\big{)}$ the measure $\mu_{\beta}$ is transformed into
@@
Thus, it remains to show that the measure $\widetilde{\mu}_{\beta}$ is invariant under the transformed evolution equation for $(z,\zeta)$. Using the linear structure of $\widetilde{M}_{t}$ we see that $(z_{t},\eta_{t})$ is a solution of (2.9) if and only if $(z_{z},\zeta_{t})=\widetilde{M}_{t}(z_{t},\eta_{t})$ solves
@@ 3.1 Compressions and dilations
While for coarse-graining compressions are the natural concept, the theory of dilations provides crucial motivation for the choice of the model discussed in this paper. For a given strongly continuous contraction semigroup (thought of as a large-scale, thermodynamic description), the theory of dilations provides the existence of a unitary semigroup which is unique up to isomorphism, and has a representation as time shift in a suitable space (the theory of dilations is closely related to one-dimensional scattering theory, cf. [LaS89, Cha. III.2]). We give an explicit construction of this dilation in Appendix LABEL:sec:Appendix-Dilations.
An important aspect of this specific construction is that it can simulaneously be used to define the larger space ${\mathbf{X}}$ and the extension of the unitary group $(\mathrm{e}^{t{\mathbb{J}}_{\mathcal{B}}})_{t\in{\mathbb{R}}}$ on ${\mathbf{H}}$ to a strongly continuous group $(\widehat{S}_{t})_{t\in{\mathbb{R}}}$ on the space ${\mathbf{X}}$, as required by Assumption 2.3.
@@ 3.2 Derivation of the coarse-grained equations
    $\nu_{\beta}({\mathrm{d}}z\;\!\mathrm{d}w)=\frac{1}{Z_{\beta}}\exp\Big{[}-\beta\Big{(}{\mathcal{H}}_{\mathcal{A}}(z)+\tfrac{1}{2}\|w\|_{{\mathbf{H}}}^{2}+\left({\mathsf{C}}z\middle|w\right)_{\mathbf{H}}^{2}\Big{)}\Big{]}\;\!\mathrm{d}z\;\!\mathrm{d}w.$ (3.5)
@@ 4.1 Deterministic GENERIC
A pre-form of this structure was introduced by Morrison [Mor86] under the name metriplectic systems, and further developed by Öttinger and Grmela as General Equation for the Non-Equilibrium Reversible-Irreversible Coupling, abbreviated as GENERIC [GrÖ97, ÖtG97, Ö05]. In [Mor86] the structures ${\mathbb{J}}$ and ${\mathbb{K}}$ feature together with one function ${\mathcal{F}}={\mathcal{E}}-{\mathcal{S}}$. The importance of keeping ${\mathcal{E}}$ and ${\mathcal{S}}$ separate and the role of the non-interaction conditions was emphasized for the first time in [GrÖ97, ÖtG97]. Grmela and Öttinger’s treatment includes extensions to stochastic evolutions, which are central to this paper, and we therefore follow the GENERIC terminology.
@@
The GENERIC equation (4.1) is coordinate-invariant: under linear or nonlinear changes of variables the equations retain the same form (see e.g. [Ö05, Sec. 1.2.4] or [Mie11, Sec. 2]).
@@ 4.2 GENERIC SDEs
Fluctuation-dissipation relations of the form (4.5) have a long history; see e.g. the discussion in [ÖPM21, MPÖ21]. The GENERIC SDE (4.4) and the fluctuation-dissipation relation (4.5) seem to appear for the first time in [GrÖ97]; Öttinger refers to equations of the form (4.4) as ‘GENERIC with fluctuations’ [Ö05, Sec. 1.2.5].
@@
To show the conservation of ${\mathcal{E}}$, we calculate by Itō’s lemma that
@@
In fact, the conservation of ${\mathcal{E}}$ by the SDE (4.4) implies that there is a very large class of invariant meaures. For instance, any measure of the form $f({\mathcal{E}}(y))\mathrm{e}^{{\mathcal{E}}(y)}\;\!\mathrm{d}y$ also is conserved, for any measurable function $f\colon{\mathbb{R}}\to{\mathbb{R}}$. As another example, in Section 4.5 we construct invariant measures $\nu_{\beta}^{{\mathcal{E}}_{0}}$ that can be interpreted as “$\mathrm{e}^{\mathcal{S}}(y)\;\!\mathrm{d}y$ restricted to a level set ${\mathcal{E}}={\mathcal{E}}_{0}$”, which can be seen as an extreme version of $f({\mathcal{E}}(y))\mathrm{e}^{{\mathcal{E}}(y)}\;\!\mathrm{d}y$.
@@
This observation suggests the following interpretation of the interplay between $m$, ${\mathcal{S}}$, and the Lebesgue measure. The Poisson structure ${\mathbb{J}}$ admits an invariant measure $\mu=m\;\!\mathrm{d}y$; often $\mu$ coincides with the Lebesgue measure. The entropy ${\mathcal{S}}$ has the interpretation of changing the weight of different $y$-states, thus modifying $\mu$ into $\mathrm{e}^{{\mathcal{S}}}\mu$ (see also Section 4.5). The gradient term ${\mathbb{K}}{\mathrm{D}}{\mathcal{S}}$ and the Itō correction term $m^{-1}\mathop{\mathrm{div}}\nolimits\left(m{\mathbb{K}}\right)$ combine to make the measure $\mathrm{e}^{{\mathcal{S}}}\mu$ invariant.
@@
See also [Ö05, (6.76) and (6.163)] for a different approach to this question.
@@
We now generalise this concept of reversibility by lifting it to measures on path space. A measure $P$ on the path space ${\mathrm{C}}((-\infty,\infty);{\mathbb{R}}^{d})$ is defined to be $R$-reversible if it is invariant under the combined action of $R$ and time inversion:
@@ 4.3 Reformulation in GENERIC form
This split suggests to introduce a new scalar variable $e=e(t)\in{\mathbb{R}}$ that keeps track of the energy that is exchanged between the macroscopic part and the invisible part of the heat bath:
@@
Using Itō’s lemma to express the right-hand side in (4.14) leads to the following system of equations in the three unknowns $(z,w,e)$. This is one of the main results of this paper.
@@
${\mathcal{E}}_{\mathrm{GEN}}(z,w,e)$ $={\mathcal{H}}_{\mathcal{A}}(z)+\frac{1}{2}\|w\|_{\mathbf{H}}^{2}+\left({\mathsf{C}}z\middle|w\right)_{\mathbf{H}}+e,$ (4.20a)
@@
For the model in Examples (1.1) and (1.3), where we identify ${\mathbf{W}}$ with ${\mathbb{R}}^{3}$ we obtain a even more specific form of the GENERIC Poisson and Onsager structure in in ${\mathbb{R}}^{(2n+4){\times}(2n+4)}$:
${\mathbb{J}}_{\mathrm{GEN}}=\left(\begin{array}[]{cccc}0&{\mathbb{I}}&0&0\\ \!\!{-}{\mathbb{I}}&0&0&0\\ 0&0&\!\!-{{\mathbb{D}}_{\mathrm{skw}}}&0\\ 0&0&0&0\end{array}\right)\ \text{ and}\quad{\mathbb{K}}_{\mathrm{GEN}}=\frac{1}{\beta}\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&{{\mathbb{D}}_{\mathrm{sym}}}&\!\!\!-{{\mathbb{D}}_{\mathrm{sym}}}(w{+}\mathsf{C}q)\!\!\\ 0&0&\!\!{-}\big{(}{{\mathbb{D}}_{\mathrm{sym}}}(w{+}\mathsf{C}q)\big{)}^{\!\top}\!\!\!&|w{+}\mathsf{C}q|^{2}_{{{\mathbb{D}}_{\mathrm{sym}}}}\end{array}\right)$
where $\mathsf{C}\in{\mathbb{R}}^{3{\times}n}$ and ${{\mathbb{D}}_{\mathrm{sym}}},\,{{\mathbb{D}}_{\mathrm{skw}}}\in{\mathbb{R}}^{3{\times}3}$ have the explicit form
@@
$\frac{{\mathrm{d}}}{{\mathrm{d}}t}\left(\begin{array}[]{c}q\\ p\\ w\\ e\end{array}\right)=\left(\begin{array}[]{c}p\\ \!\!-{\mathrm{D}}V(q)-\mathsf{C}^{*}w\!\\ -{{\mathbb{D}}_{\mathrm{skw}}}(w{+}\mathsf{C}q)\\ 0\end{array}\right)+\left(\begin{array}[]{c}0\\ 0\\ \!\!{-}{{\mathbb{D}}_{\mathrm{sym}}}(w{+}\mathsf{C}q)\!\\ |w{+}\mathsf{C}q|^{2}_{{{\mathbb{D}}_{\mathrm{sym}}}}\end{array}\right)=\left(\begin{array}[]{c}p\\ -{\mathrm{D}}V(q)-\mathsf{C}^{*}w\\ -{\mathbb{D}}(w{+}\mathsf{C}q)\\ |w{+}\mathsf{C}q|^{2}_{{{\mathbb{D}}_{\mathrm{sym}}}}\end{array}\right).$
@@
It would be interesting to understand in what sense or in which cases the formula (4.20g) for ${\mathbb{J}}_{\mathrm{GEN}}$ coincides with Öttinger’s coarse-graining formula in [Ö05, Eq. (6.67+68)]:
@@
Thus this formula should be considered in a renormalized sense, i.e. be approximating ${\mathbb{J}}_{\mathcal{B}}$ suitably. Because of the generator property, we have the approximation property by a symmetric difference of going forward and backward in time (which means extracting irreversible information):
@@
$\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|w_{t}\|_{\mathbf{H}}^{2}=\left(w_{t}\middle|{\mathbb{D}}w_{t}\right)_{\mathbf{H}}+\left({\mathbb{D}}w_{t}\middle|w_{t}\right)_{\mathbf{H}}=2\left(w_{t}\middle|{{\mathbb{D}}_{\mathrm{sym}}}w_{t}\right)_{\mathbf{H}}.$

### v3

@@ 2.4 Invariant measure on the product space ${\mathbf{Z}}{\times}{\mathbf{X}}$
$\exp\left(-\beta{\mathcal{H}}_{\mathrm{total}}(z,\eta)\right)=\exp\left(-\beta\big{(}{\mathcal{H}}_{\mathcal{A}}(z)-\frac{1}{2}\|{\mathsf{C}}z\|_{\mathbf{H}}^{2}\big{)}\right)\exp\left(-\frac{\beta}{2}\|\eta{+}{\mathsf{C}}z\|_{\mathbf{H}}^{2}\right).$
@@
By introducing the new variables $(z_{t},\zeta_{t})=\widetilde{M}_{t}(z,\eta)=\big{(}z,\widehat{S}_{-t}(\eta{+}{\mathsf{C}}z)\big{)}$ the measure $\mu_{\beta}$ is transformed into
@@
Thus, it remains to show that the measure $\widetilde{\mu}_{\beta}$ is invariant under the transformed evolution equation for $(z,\zeta)$. Using the linear structure of $\widetilde{M}_{t}$ we see that $(z_{t},\eta_{t})$ is a solution of (2.9) if and only if $(z_{t},\zeta_{t})=\widetilde{M}_{t}(z_{t},\eta_{t})$ solves
@@ 3.1 Compressions and dilations
While for coarse-graining compressions are the natural concept, the theory of dilations provides crucial motivation for the choice of the model discussed in this paper. For a given strongly continuous contraction semigroup (thought of as a large-scale, thermodynamic description), the theory of dilations provides the existence of a unitary semigroup which is unique up to isomorphism, and has a representation as time shift in a suitable space (the theory of dilations is closely related to one-dimensional scattering theory, cf. [LaP89, Cha. III.2]). We give an explicit construction of this dilation in Appendix LABEL:sec:Appendix-Dilations.
An important aspect of this specific construction is that it can simultaneously be used to define the larger space ${\mathbf{X}}$ and the extension of the unitary group $(\mathrm{e}^{t{\mathbb{J}}_{\mathcal{B}}})_{t\in{\mathbb{R}}}$ on ${\mathbf{H}}$ to a strongly continuous group $(\widehat{S}_{t})_{t\in{\mathbb{R}}}$ on the space ${\mathbf{X}}$, as required by Assumption 2.3.
@@ 3.2 Derivation of the coarse-grained equations
    $\nu_{\beta}({\mathrm{d}}z\;\!\mathrm{d}w)=\frac{1}{Z_{\beta}}\exp\Big{[}-\beta\Big{(}{\mathcal{H}}_{\mathcal{A}}(z)+\tfrac{1}{2}\|w\|_{{\mathbf{H}}}^{2}+\left({\mathsf{C}}z\middle|w\right)_{\mathbf{H}}\Big{)}\Big{]}\;\!\mathrm{d}z\;\!\mathrm{d}w.$ (3.5)
@@ 4.1 Deterministic GENERIC
A pre-form of this structure was introduced by Grmela in [Grm84, Grm85] and independently by Morrison [Mor84, Mor86] under the name metriplectic systems, and further developed by Öttinger and Grmela as General Equation for the Non-Equilibrium Reversible-Irreversible Coupling, abbreviated as GENERIC [GrÖ97, ÖtG97, Ött05]. In [Mor86] the structures ${\mathbb{J}}$ and ${\mathbb{K}}$ feature together with one function ${\mathcal{F}}={\mathcal{E}}-{\mathcal{S}}$. The importance of keeping ${\mathcal{E}}$ and ${\mathcal{S}}$ separate and the role of the non-interaction conditions was emphasized for the first time in [GrÖ97, ÖtG97]. Grmela and Öttinger’s treatment includes extensions to stochastic evolutions, which are central to this paper, and we therefore follow the GENERIC terminology.
@@
The GENERIC equation (4.1) is coordinate-invariant: under linear or nonlinear changes of variables the equations retain the same form (see e.g. [Ött05, Sec. 1.2.4] or [Mie11, Sec. 2]).
@@ 4.2 GENERIC SDEs
Fluctuation-dissipation relations of the form (4.5) have a long history; see e.g. the discussion in [ÖPM21, MPÖ21]. The GENERIC SDE (4.4) and the fluctuation-dissipation relation (4.5) seem to appear for the first time in [GrÖ97]; Öttinger refers to equations of the form (4.4) as ‘GENERIC with fluctuations’ [Ött05, Sec. 1.2.5].
@@
To show the conservation of ${\mathcal{E}}$, we calculate by Itô’s lemma that
@@
In fact, the conservation of ${\mathcal{E}}$ by the SDE (4.4) implies that there is a very large class of invariant measures. For instance, any measure of the form $f({\mathcal{E}}(y))\mathrm{e}^{{\mathcal{E}}(y)}\;\!\mathrm{d}y$ also is conserved, for any measurable function $f\colon{\mathbb{R}}\to{\mathbb{R}}$. As another example, in Section 4.5 we construct invariant measures $\nu_{\beta}^{{\mathcal{E}}_{0}}$ that can be interpreted as “$\mathrm{e}^{\mathcal{S}}(y)\;\!\mathrm{d}y$ restricted to a level set ${\mathcal{E}}={\mathcal{E}}_{0}$”, which can be seen as an extreme version of $f({\mathcal{E}}(y))\mathrm{e}^{{\mathcal{E}}(y)}\;\!\mathrm{d}y$.
@@
This observation suggests the following interpretation of the interplay between $m$, ${\mathcal{S}}$, and the Lebesgue measure. The Poisson structure ${\mathbb{J}}$ admits an invariant measure $\mu=m\;\!\mathrm{d}y$; often $\mu$ coincides with the Lebesgue measure. The entropy ${\mathcal{S}}$ has the interpretation of changing the weight of different $y$-states, thus modifying $\mu$ into $\mathrm{e}^{{\mathcal{S}}}\mu$ (see also Section 4.5). The gradient term ${\mathbb{K}}{\mathrm{D}}{\mathcal{S}}$ and the Itô correction term $m^{-1}\mathop{\mathrm{div}}\nolimits\left(m{\mathbb{K}}\right)$ combine to make the measure $\mathrm{e}^{{\mathcal{S}}}\mu$ invariant.
@@
See also [Ött05, (6.76) and (6.163)] for a different approach to this question.
@@
We now generalize this concept of reversibility by lifting it to measures on path space. A measure $P$ on the path space ${\mathrm{C}}((-\infty,\infty);{\mathbb{R}}^{d})$ is defined to be $R$-reversible if it is invariant under the combined action of $R$ and time inversion:
@@ 4.3 Reformulation in GENERIC form
This split suggests introducing a new scalar variable $e=e(t)\in{\mathbb{R}}$ that keeps track of the energy that is exchanged between the macroscopic part and the invisible part of the heat bath:
@@
Using Itô’s lemma to express the right-hand side in (4.14) leads to the following system of equations in the three unknowns $(z,w,e)$. This is one of the main results of this paper.
@@
${\mathcal{E}}_{\mathrm{GEN}}(z,w,e)$ $={\mathcal{H}}_{\mathcal{A}}(z)+\frac{1}{2}\|w\|_{\mathbf{H}}^{2}+\left({\mathsf{C}}z\middle|w\right)_{\mathbf{H}}+e$ (4.20a)
@@
For the model in Examples (1.1) and (1.4), where we identify ${\mathbf{W}}$ with ${\mathbb{R}}^{3}$ we obtain an even more specific form of the GENERIC Poisson and Onsager structure in ${\mathbb{R}}^{(2n+4){\times}(2n+4)}$:
${\mathbb{J}}_{\mathrm{GEN}}=\left(\begin{array}[]{cccc}0&{\mathbb{I}}&0&0\\ \!\!{-}{\mathbb{I}}&0&0&0\\ 0&0&\!\!-{{\mathbb{D}}_{\mathrm{skw}}}&0\\ 0&0&0&0\end{array}\right)\ \text{ and}\quad{\mathbb{K}}_{\mathrm{GEN}}=\frac{1}{\beta}\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&0&0\\ 0&0&{{\mathbb{D}}_{\mathrm{sym}}}&\!\!\!-{{\mathbb{D}}_{\mathrm{sym}}}(w{+}{\mathsf{K}}q)\!\!\\ 0&0&\!\!{-}\big{(}{{\mathbb{D}}_{\mathrm{sym}}}(w{+}{\mathsf{K}}q)\big{)}^{\!\top}\!\!\!&|w{+}{\mathsf{K}}q|^{2}_{{{\mathbb{D}}_{\mathrm{sym}}}}\end{array}\right)$
where ${\mathsf{K}}\in{\mathbb{R}}^{3{\times}n}$ and ${{\mathbb{D}}_{\mathrm{sym}}},\,{{\mathbb{D}}_{\mathrm{skw}}}\in{\mathbb{R}}^{3{\times}3}$ have the explicit form
@@
$\frac{{\mathrm{d}}}{{\mathrm{d}}t}\left(\begin{array}[]{c}q\\ p\\ w\\ e\end{array}\right)=\left(\begin{array}[]{c}p\\ \!\!-{\mathrm{D}}V(q)-{\mathsf{K}}^{*}w\!\\ -{{\mathbb{D}}_{\mathrm{skw}}}(w{+}{\mathsf{K}}q)\\ 0\end{array}\right)+\left(\begin{array}[]{c}0\\ 0\\ \!\!{-}{{\mathbb{D}}_{\mathrm{sym}}}(w{+}{\mathsf{K}}q)\!\\ |w{+}{\mathsf{K}}q|^{2}_{{{\mathbb{D}}_{\mathrm{sym}}}}\end{array}\right)=\left(\begin{array}[]{c}p\\ -{\mathrm{D}}V(q)-{\mathsf{K}}^{*}w\\ -{\mathbb{D}}(w{+}{\mathsf{K}}q)\\ |w{+}{\mathsf{K}}q|^{2}_{{{\mathbb{D}}_{\mathrm{sym}}}}\end{array}\right).$
@@
It would be interesting to understand in what sense or in which cases the formula (4.20g) for ${\mathbb{J}}_{\mathrm{GEN}}$ coincides with Öttinger’s coarse-graining formula in [Ött05, Eq. (6.67+68)]:
@@
Thus this formula should be considered in a renormalized sense, i.e. by approximating ${\mathbb{J}}_{\mathcal{B}}$ suitably. Because of the generator property, we have the approximation property by a symmetric difference of going forward and backward in time (which means extracting irreversible information):
@@
$\frac{{\mathrm{d}}}{{\mathrm{d}}t}\|w_{t}\|_{\mathbf{H}}^{2}=-\left(w_{t}\middle|{\mathbb{D}}w_{t}\right)_{\mathbf{H}}-\left({\mathbb{D}}w_{t}\middle|w_{t}\right)_{\mathbf{H}}=-2\left(w_{t}\middle|{{\mathbb{D}}_{\mathrm{sym}}}w_{t}\right)_{\mathbf{H}}.$
@@ 4.6 Characterization of entropy through finite-dimensional approximations
4.7 Reduction to a dissipative system A via strong coupling and fast relaxation
In applications it is often interesting to study a damped version of the Hamiltonian system A coupled directly to the heat bath, such that no additional internal variable $w\in{\mathbf{W}}$ appears. We discuss here on the formal level how such models can occur as a limit of systems including $w$. The idea is to make the coupling between $w$ and $z$ very strong by setting ${\mathsf{C}}=\lambda\widetilde{\mathsf{C}}$ and to make the damping matrix even stronger by setting ${\mathbb{D}}=\lambda^{2}\widetilde{\mathbb{D}}$. Here $\lambda\gg 1$ is a large real parameter such that $\lambda^{2}$ measures the quotient of the slow time scales in System A and the fast relaxation time for $w$. We will be interested in the limit $\lambda\to\infty$ while keeping the (inverse) temperature $\beta>0$ fixed.
We start from the SDE (4.15) for the variables $(z_{t},w_{t},e_{t})$ and assume that ${\mathcal{H}}_{\mathcal{A}}$ is given in the form ${\mathcal{H}}_{\mathcal{A}}(z)=\widetilde{\mathcal{H}}(z)+\frac{1}{2}\lambda^{2}\|\widetilde{\mathsf{C}}z\|^{2}_{\mathbf{H}}$ (cf. (2.3)). For simplicity, we omit the equation for the energy $e_{t}$, and obtain
$\binom{{\mathrm{d}}z_{t}}{{\mathrm{d}}w_{t}}=\binom{{\mathbb{J}}_{\mathcal{A}}\big{(}{\mathrm{D}}\widetilde{\mathcal{H}}(z_{t})+\lambda\widetilde{\mathsf{C}}^{*}(w_{t}{+}\lambda\widetilde{\mathsf{C}}z_{t})\big{)}}{-\lambda^{2}\widetilde{\mathbb{D}}(w_{t}{+}\lambda\widetilde{\mathsf{C}}z_{t})}\;\!\mathrm{d}t+\binom{0}{\lambda\widetilde{\Sigma}{\mathrm{d}}B_{t}}\text{ with }\widetilde{\Sigma}\widetilde{\Sigma}^{*}=\frac{1}{\beta}(\widetilde{\mathbb{D}}{+}\widetilde{\mathbb{D}}^{*}).$ (4.66)
The equation for $w$ is linear and it follows that for $\lambda\gg 1$ one needs $w_{t}\approx-\lambda{\mathsf{C}}z_{t}$ to cancel the highest order terms in $\lambda$. To see the influence of $w_{t}$ on the macroscopic variable $z_{t}$, it is advantageous to introduce the new variable $u^{\lambda}=\lambda(w{+}\lambda{\mathsf{C}}z)$, which satisfies the SDE
${\mathrm{d}}z_{t}$ $={\mathbb{J}}_{\mathcal{A}}(z_{z}){\mathrm{D}}\widetilde{\mathcal{H}}(z_{t})+{\mathbb{J}}_{\mathcal{A}}(z_{z})\widetilde{\mathsf{C}}^{*}\;\!\mathrm{d}u^{\lambda}_{t},$
${\mathrm{d}}u^{\lambda}_{t}$ $=-\lambda^{2}\big{(}\,\widetilde{\mathbb{D}}-\widetilde{\mathsf{C}}{\mathbb{J}}_{\mathcal{A}}(z_{t})\widetilde{\mathsf{C}}^{*}\big{)}u^{\lambda}_{t}\;\!\mathrm{d}t+\lambda^{2}\widetilde{\mathsf{C}}{\mathbb{J}}_{\mathcal{A}}(z_{t}){\mathrm{D}}\widetilde{\mathcal{H}}(z_{t})+\lambda^{2}\widetilde{\Sigma}\;\!\mathrm{d}B_{t}.$
For simplicity, we now further assume that ${\mathbb{J}}_{\mathcal{A}}$ is independent of $z$, so that $\widetilde{\mathbb{G}}:=\widetilde{\mathbb{D}}-\widetilde{\mathsf{C}}{\mathbb{J}}_{\mathcal{A}}\widetilde{\mathsf{C}}^{*}$ satisfies $\widetilde{\mathbb{G}}{+}\widetilde{\mathbb{G}}^{*}=\widetilde{\mathbb{D}}{+}\widetilde{\mathbb{D}}^{*}>0$ and hence generates a contraction semigroup $\big{(}\mathrm{e}^{-t\widetilde{\mathbb{G}}}\big{)}_{t\geq 0}$. With this, $u^{\lambda}_{t}$ can be expressed by Duhamel’s formula as
$u^{\lambda}_{t}=\mathrm{e}^{-t\lambda^{2}\widetilde{\mathbb{G}}}u_{0}+\int_{0}^{t}\lambda^{2}\,\mathrm{e}^{-(t-s)\lambda^{2}\widetilde{\mathbb{G}}}\>\widetilde{\mathsf{C}}{\mathbb{J}}_{\mathcal{A}}{\mathrm{D}}\widetilde{\mathcal{H}}(z_{s})\;\!\mathrm{d}s+\int_{0}^{t}\lambda^{2}\,\mathrm{e}^{-(t-s)\lambda^{2}\widetilde{\mathbb{G}}}\>\widetilde{\Sigma}\;\!\mathrm{d}B_{s}.$
From this, we find the deterministic limit $U_{t}=\lim_{\lambda\to\infty}u^{\lambda}_{t}=\widetilde{\mathbb{G}}^{-1}\widetilde{\mathsf{C}}{\mathbb{J}}_{\mathcal{A}}{\mathrm{D}}\widetilde{\mathcal{H}}(z_{t})+\widetilde{\mathbb{G}}^{-1}\widetilde{\Sigma}\,B_{t}$, and inserting this relation into the equation for $z_{t}$ yields an SDE for $z_{t}$ alone, namely
$\;\!\mathrm{d}z_{t}={\mathbb{J}}_{\mathcal{A}}{\mathrm{D}}\widetilde{\mathcal{H}}(z_{t})\;\!\mathrm{d}t+{\mathbb{J}}_{\mathcal{A}}\widetilde{\mathsf{C}}^{*}\;\!\mathrm{d}U_{t}=\big{(}\,\widehat{\mathsf{J}}-\widehat{\mathsf{D}}\big{)}{\mathrm{D}}\widetilde{\mathcal{H}}(z_{t})\;\!\mathrm{d}t+\widehat{\bm{\Sigma}}\;\!\mathrm{d}B_{t}$ (4.67)
with $\text{with }\widehat{\mathsf{J}}={\mathbb{J}}_{\mathcal{A}}-\tfrac{1}{2}{\mathbb{J}}_{\mathcal{A}}\widetilde{\mathsf{C}}^{*}(\widetilde{\mathbb{G}}^{-1}{-}\widetilde{\mathbb{G}}^{-*})\widetilde{\mathsf{C}}\,{\mathbb{J}}_{\mathcal{A}}^{*},$
$\widehat{\mathsf{D}}=\tfrac{1}{2}{\mathbb{J}}_{\mathcal{A}}\widetilde{\mathsf{C}}^{*}(\widetilde{\mathbb{G}}^{-1}{+}\widetilde{\mathbb{G}}^{-*})\widetilde{\mathsf{C}}\,{\mathbb{J}}_{\mathcal{A}}^{*},\text{ and }\widehat{\bm{\Sigma}}={\mathbb{J}}_{\mathcal{A}}\widetilde{\mathsf{C}}^{*}\widetilde{\mathbb{G}}^{-1}\widetilde{\Sigma}.$
This SDE can be seen as the limit for $\lambda\to\infty$ of the scaled SDE (4.66). Clearly, $\widehat{\mathsf{J}}=-\widehat{\mathsf{J}}^{*}$ and $\widehat{\mathsf{D}}=\widehat{\mathsf{D}}^{*}\geq 0$ (since $\widetilde{\mathbb{G}}$ has positive symmetric part), hence (4.67) is a damped Hamiltonian system with stochastic forcing.
It remains unclear whether in the general case the SDE (4.67) satisfies the fluctuation-dissipation relation and has a simple invariant measure. However, for the special case $\widetilde{\mathsf{C}}{\mathbb{J}}_{\mathcal{A}}\widetilde{\mathsf{C}}^{*}=0$ we have $\widetilde{\mathbb{G}}=\widetilde{\mathbb{D}}$, and from $\widetilde{\mathbb{D}}^{-1}\widetilde{\Sigma}\widetilde{\Sigma}^{*}\widetilde{\mathbb{D}}^{-*}=\frac{1}{\beta}(\widetilde{\mathbb{D}}^{-1}{+}\widetilde{\mathbb{D}}^{-*})$ the fluctuation-dissipation relation
$\widehat{\bm{\Sigma}}\widehat{\bm{\Sigma}}^{*}=\frac{1}{\beta}\,{\mathbb{J}}_{\mathcal{A}}\widetilde{\mathsf{C}}^{*}\big{(}\widetilde{\mathbb{D}}^{-1}{+}\widetilde{\mathbb{D}}^{-*}\big{)}\widetilde{\mathsf{C}}{\mathbb{J}}_{\mathcal{A}}^{*}=\frac{2}{\beta}\widehat{\mathsf{D}}$
follows. With this we see that (4.67) has the invariant measure $\widehat{\nu}_{\beta}({\mathrm{d}}z)=Z_{\beta}^{-1}\mathrm{e}^{-\beta\widetilde{H}(z)}\;\!\mathrm{d}z$, cf. (3.5).
Example 4.16 (Molecular dynamics).
[…]
$\widetilde{\mathsf{C}}{\mathbb{J}}\widetilde{\mathsf{C}}^{*}=0,\quad\widetilde{\mathbb{G}}=\widetilde{\mathbb{D}},\quad\widehat{\mathsf{D}}={\mathsf{K}}^{*}\widetilde{\mathbb{D}}^{-1}{\mathsf{K}},\quad\widehat{\bm{\Sigma}}=\binom{0}{-\sqrt{2/\beta}\>{\mathsf{K}}^{*}\widetilde{\mathbb{D}}^{-1/2}}.$
Hence, the fluctuation-dissipation relation holds, and the classical invariant measure reads
$\widehat{\nu}_{\beta}({\mathrm{d}}z)=\frac{1}{Z_{\beta}}\>\mathrm{e}^{-\beta\widetilde{\mathcal{H}}(z)}\;\!\mathrm{d}z=\frac{1}{Z_{\beta}}\,\exp\Big{(}\!-\beta\Phi(z)-\frac{\beta}{2}p^{*}M^{-1}p\Big{)},$
i.e. the momentum has a Gaussian (or Maxwellian) distribution.